A company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 99.5-cm and a standard deviation of 2.5-cm.
Find the probability that the length of a randomly selected steel rod is between 99-cm and 106-cm.
P(99-cm < X < 106-cm) =
I tried doing the length times the standard deviation.

Answer by stanbon(75887)

A company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 99.5-cm and a standard deviation of 2.5-cm.
Find the probability that the length of a randomly selected steel rod is between 99-cm and 106-cm.
----
z(99) = (99-99.5)/2.5 = 0.5/2.5 = 1/5
---
z(106) = (106-99.5)/2.5 = 2.6
----
P(99< x <106) = P(1/5< z < 2.6) = normalcdf(1/5,2.6) = 0.4161
=====================
Cheers,
Stan H.
=======================